# Navigating the Lobbying Landscape: Insights from Opinion Dynamics Models

### v2

@@ Introduction
In all simulations, we monitor various criteria. An important aspect is the number of clusters that form in the population: consensus means one cluster, more clusters mean fragmentation and polarisation. We employ the effective number of clusters defined as $C={N^{2}}/{\sum_{i}^{k}N_{i}^{2}}$, where $k$ is the number of clusters and $N_{i}$ is the size of each cluster. This measure takes into account not only the number of groups, but also their size. $C$ is equal to $k$ when we have $k$ clusters of equal size, and it is smaller when the clusters are imbalanced. A second criterion of interest is the average opinion of agents at the end of the simulation, which we report in terms of average of the $p_{i,\tau}$ values as defined in Eq. 1. In addition to indicating how beliefs have collectively shifted, this criterion is also informative of lobbyists’ payoff. Furthermore, we look at the number of iterations required for convergence, indicating the speed of opinion cluster formation. For all parameter configurations we provide mean values of these criteria over 150 independent runs, each one of which is simulated until the model reaches convergence (i.e., until all the individual weights assigned to the competing models stop updating).
@@ Results
(b) Number of clusters: zoom on the bottom right corner of the parameter space
@@
(d) Average subjective probabilities: zoom on the bottom right corner of the parameter space
@@
In the first setting, we investigated the properties of the model in the absence of any active lobbyist, when the nodes are left free to interact with their peers without outside input. Figure 1 displays the effective number of clusters and the average subjective probabilities across the parameter space when the stable state is reached. We observe that, for most of the parameter region under consideration, the distribution of subjective probabilities tends to converge to a general consensus, with all of the nodes eventually supporting a single model (either $o$ or $p$, depending on initial conditions). This dynamic is relatively quick and is also illustrated by the evolution plot in Fig. 2(a).
@@
Two notable exceptions to this behavior are worth mentioning: the first, and more trivial, occurs in the special case where $\lambda=1$ and $\phi=0$. In this case, as it is immediately clear from equation (3), nodes are “deaf” to whatever signal they receive from their peers, and any update of subjective probabilities is impossible. In this case, there is no dynamics and each agent maintains its initial condition. The second exception emerges at the bottom and to the right of the parametric region, where the confirmation bias parameter $\phi$ is sufficiently large to prevent the network from reaching a consensus. In this case, in many runs the model converges to a polarized equilibrium, in which a variable proportion of agents supports one model, while the rest of the network supports the competing one (see Fig. 2(b) for an example). This indicates that confirmation bias has an important role in our model, facilitating polarization of opinions, since agents are influenced little by peers with conflicting opinions. The effect is larger as lambda increases, i.e. as under-reaction is growing, confirmation bias becomes more effective in generating polarization. This seems to indicate that a society that changes opinion slowly is more susceptible to polarisation due to confirmation bias.
@@
Another basic property of the model is that, while the model reaches consensus generally very fast, the time required to reach the stable state is increasing in $\lambda$ (as shown in Fig. 3). In particular, as $\lambda$ approaches 1, nodes take a larger number of iterations to reach a stable configuration, because each of them is less sensitive to new information and under-reacts. The relation between the convergence time and $\phi$ is, however, not so linear. For low $\lambda$ values, large $\phi$ seems to slow down convergence, i.e. if the individuals do not under-react, then confirmation bias slows down convergence as society polarises. However, when $\lambda$ is large, the opposite effect is visible: since individuals under-react, polarisation is faster as confirmation bias increases: they are not attracted by peers from the opposite opinion groups and so the simulations ends faster.
@@
In Fig. 4 we show the number of clusters and average subjective probabilities at the steady state. While in the baseline simulation outcomes were balanced between the two competing models (thus resulting in average subjective probabilities around 0.5 across independent runs, see Fig. 1), in the one-lobbyist scenario average probabilities are significantly higher over the whole parameter space, indicating that generally the lobbyist is able to steer part of the network towards supporting its own preferred model. Furthermore, the effectiveness of the lobbyist is not globally homogeneous over the parameter space: in fact, the average probabilities are around $0.99$ in the bottom-left corner region (indicating that, with these configurations, the lobbyist is successful in influencing the network almost always), but tend to be lower as we go farther off said region. Moreover, we see that the lobbyist, while managing to successfully increase the average subjective probability distributions, is unable to eliminate or reduce clustering in the bottom-right corner of the map (Fig. 4). If anything, in this region of the map an increase in polarization emerges, as the effective number of clusters increases, i.e. the size of the second cluster increases.
@@
These results indicate that a society with medium-high levels of under-reaction and low confirmation bias (dark blue area in Fig. 4(b)) are easily influenced by a lobbyist, while when confirmation bias is strong or the population is more dynamic (low under-reaction), the peer-effect can contrast the lobbyist. This can generate consensus on the opposite opinion in a minority of simulations, when confirmation bias is low, or increase polarisation when confirmation bias is high. There appears to be a type of phase transition at the boundary, where the lobbyist does not have full power any more.
@@
Figure 4: One-lobbyist scenario. The effective number of clusters and average subjective probability of the final opinion distribution is represented as a function of the strength of directional motivated reason $\phi$ and the degree of under-reaction $\lambda$ in presence of a lobbyist. The lobbyists supports the pessimistic model, has a budget $B=10,\!000$ to send its signals and can be active for a time horizon $T=100$.
To investigate further this phase transition, we performed budget sensitivity analysis, i.e. changed the budget of the lobbyist from 10,000 to 5,000, 20,000 and 40,000, allowing them to reach 10%, 40% and 80% of the population, respectively. In Fig. 5, we show the resulting average opinions. It is very clear that the region in which the lobbyist dominates the network (where it can impose its model almost certainly) expands linearly with the amount of resources available. With $B>40,\!000$, the lobbyist influence is complete and global over the parameter space. Even outside said region, however, average subjective probabilities are higher on average the higher the budget allocated to the lobbyist, indicating that the lobbyist has still a non-negligible impact on the network behaviour regardless of the specific parameter configuration.
(a) $B=5,\!000$
(b) $B=10,\!000$
(c) $B=20,\!000$
(d) $B=40,\!000$
@@
In this configuration, average subjective probabilities across simulations seem to balance out, just as in the baseline scenario, as it is shown in Fig. 6(b). In fact, as the two lobbyists are identical except for the specific model supported, their competing influence seems to cancel out, and the average subjective probabilities gravitates around $\bar{p}=0.5$, regardless of the parametrization.
@@
Figure 6: The two-lobbyists scenario. The average number of iterations and average subjective probabilities of the final opinion distribution are represented as a function of strength of directional motivated reason $\phi$ and the degree of under-reaction $\lambda$. The lobbyists supports opposing models (one the optimistic model and the other the pessimistic one). Both of them are endowed with a budget $B=10,\!000$ to send its signals to the agents, i.e. can reach out an average of $20\%$ of the agents in each time steps and can be active for a time horizon $T=100$.
@@
The main difference emerging in this case in comparison to the baseline scenario is that competition between lobbyists over influence on the network results in an increase in the number of iterations needed to reach the stable state, as shown in Fig. 6(a).
@@
In other words, the joint activity of the two lobbyists prevents the network from reaching an equilibrium as long as the lobbyists send signals; this is particularly true in the region that, in the previous scenario, was associated with higher lobbyist efficacy, strengthening the impression that, for that subset of configurations, it is easier for the external agent to project its influence on the network. Figure 6(c) represents the evolution of a typical run of the model in one of those configurations. Differently from what happens in the baseline scenario, where the network converges to a stable state relatively quickly, the system undergoes persistent fluctuations for as long as the two lobbyists are active on the network. Shortly after they stop emitting signals (beyond 100 iterations), however, the network quickly converges to a consensus, in a similar dynamics to what we have seen in the baseline scenario.
@@
In this last scenario, we try to go a step further and we perform a sensitivity analysis on the lobbyists’ strategic horizon $T$ with the aim of investigating how this parameter affects the model’s properties. Specifically, we run several different batches of simulations by increasing the parameter $T$ to 1,000 and 2,000. At the same time, we adjust the lobbyists’ budget $B$ accordingly, in order to maintain the same rate of signals per time-period of 20%, thus keeping the results comparable across specifications. As shown in Fig. 7, with a longer time horizon generally average opinion $\bar{p}$ remains about 0.5 in most configurations, suggesting that a longer horizon by the lobbyists does not radically change the outcome of the simulations. The average number of iterations, instead, tends to increase linearly with $T$ in the bottom-left corner of the parameter space, the region where the lobbyists are most effective , while it does not seem to change much in the other regions, that converge very fast due to peer effects.
(a) Average subjective probabilities for $T=1,\!000$
(b) Average subjective probabilities for $T=2,\!000$
(c) Zoom of average subjective probabilities for $T=1,\!000$
(d) Zoom of average subjective probabilities for $T=2,\!000$
(e) Average number of iterations for $T=1,\!000$
(f) Average number of iterations for $T=2,\!000$
@@
An important result is that, despite the complete symmetry of our model and of the two lobbyist agents, when increasing $T$ we observe the emergence of a locus of configurations in which average probabilities are markedly lower than 0.5 (see middle row panes of Fig. 7). This locus coincides with the transition between the two phases, where we switch from a prevailing lobbyist to a prevailing peer effect. This behaviour is more marked for longer $T$. In fact, for initial simulations with $T=100$ it was not present, while it is clearly evident for $T=1,\!000$, and $T=2,\!000$. Therefore, it appears that, for very specific combinations of under-reaction in the range [0.45,0.85] and low confirmation bias, which sit at the border between the two regimes of the dynamics, the optimistic model prevails in the presence of two identical lobbyists. This suggests that the model has a mechanism that pushes the simulation towards a certain outcome in these specific situations: the long oscillatory behaviour as the lobbyists are active seems to drift towards the optimistic model, which will eventually prevail in a large fraction of simulations. In the supplementary material we give examples of the dynamics (see Figure S3) and we also show that this behaviour does not depend on the size of the network (Figures S4 and S5). In order to rule out numerical bias, we tried to invert the model specifications, which resulted in an inversion of the prevailing model in the same area. Therefore we cannot ascribe this effect to numerical loss of precision, but to the model dynamics. Further analyses are required to understand whether this behaviour can be associated with some real outcome in opinion formation.
@@ Discussion
In this work we introduced a novel model of opinion formation, where opinions change through a learning process that resembles Bayesian learning and is characterised by cognitive biases, specifically under-reaction and directional motivated reasoning (confirmation bias). The model includes lobbyists that are characterised by strategies (a set of agents to target at each iteration) and are budget constrained. After having provided some game theoretic examples, we studied the model numerically and observed the emergence of two polarised clusters when confirmation bias is large, even in the absence of lobbying. One lobbyist with an uniform strategy can attract the entire population to the opinion it promotes if the budget is large enough or if the individuals have low confirmation bias and medium under-reaction levels. Otherwise, a polarised minority continues to coexist with the cluster that agrees with the lobbyist. In the presence of two opposing lobbyists, the final steady states obtained are similar to those without lobbying, just that convergence happens only after the lobbyists stop influencing the network. As long as they are active, the population continues to oscillate between different opinions without converging. This suggests that continuous lobbying can preclude consensus or convergence.
@@
Lobbyists are an important element of novelty in our model, since they are represented as more than an external field [13, 14, 15, 22, 17] or zealot agent[23]. Specifically, their effect on other agents can vary in time (at every iteration a subset of agents is reached) and is limited by a budget. This allows to model complex strategies and scenarios without changing the baseline model definition. We have shown that with this definition, one lobbyist attracts the complete population only in a certain area of the parameter space, a lobbyist influence area, where the population has low confirmation bias and medium levels of under-reaction. Polarisation appears as the population is more flexible (low under-reaction) or has more confirmation bias, where the peer effect can counteract the lobbyist influence. This polarising effect has also been seen in other models where extreme external information can cause minority extremist groups [15, 22]. A phase transition seems to appear at the boundary between these two regimes, i.e. lobbyist-influence area versus peer-effect area. When two symmetric opposing lobbyists are present, a rich behaviour is observed. During the active time horizon of the lobbyists, inside the lobbyist influence area, we observe fluctuations of opinions between the two extreme opinions, with convergence as soon as the lobbyists become silent. When oscillations are long, we observe an interesting convergence to the optimistic model for very specific parameter values at the boundary between the two regimes.
Future work will investigate more complex lobbyist strategies, extracted from real lobbying data [33]. In particular, we are interested in their effect in the two different parameter regimes. Furthermore, the phase transition will be studied in more detail, to fully understand its dynamics and, ideally, find ways to validate it with real data. Further simulations are also planned to understand the dynamics on more realistic social networks, such as scale-free graphs that better represent real societies. Moreover, integration with statistical model checking environments such as MultiVeStA [34, 35] could provide further insight into model properties and dynamics.

### v3

@@ Introduction
In all simulations, we monitor various criteria. An important aspect is the number of clusters that form in the population: consensus means one cluster, more clusters mean fragmentation and polarisation. We employ the effective number of clusters defined as $C={N^{2}}/{\sum_{i}^{k}N_{i}^{2}}$, where $k$ is the number of clusters and $N_{i}$ is the size of each cluster. This measure takes into account not only the number of groups, but also their size. $C$ is equal to $k$ when we have $k$ clusters of equal size, and it is smaller when the clusters are imbalanced. A second criterion of interest is the average opinion of agents at the end of the simulation, which we report in terms of average of the $p_{i,\tau}$ values as defined in equation (1). In addition to indicating how beliefs have collectively shifted, this criterion is also informative of lobbyists’ payoff. Furthermore, we look at the number of iterations required for convergence, indicating the speed of opinion cluster formation. For all parameter configurations we provide mean values of these criteria over 150 independent runs, each one of which is simulated until the model reaches convergence (i.e., until all the individual weights assigned to the competing models stop updating).
@@ Results
(b) Number of clusters: bottom-right corner zoom
@@
(d) Average subjective probabilities: bottom-right corner zoom
@@
In the first setting, we investigated the properties of the model in the absence of any active lobbyist, when the nodes are left free to interact with their peers without outside input. Figure 1 displays the effective number of clusters and the average subjective probabilities across the parameter space when the stable state is reached. We observe that, for most of the parameter region under consideration, the distribution of subjective probabilities tends to converge to a general consensus, with all of the nodes eventually supporting a single model (either $o$ or $p$, depending on initial conditions). This dynamic is relatively quick and is also illustrated by the evolution plot in Figure 2(a).
@@
Two notable exceptions to this behavior are worth mentioning: the first, and more trivial, occurs in the special case where $\lambda=1$ and $\phi=0$. In this case, as it is immediately clear from equation (3), nodes are “deaf” to whatever signal they receive from their peers, and any update of subjective probabilities is impossible. In this case, there is no dynamics and each agent maintains its initial condition. The second exception emerges at the bottom and to the right of the parametric region, where the confirmation bias parameter $\phi$ is sufficiently large to prevent the network from reaching a consensus. In this case, in many runs the model converges to a polarized equilibrium, in which a variable proportion of agents supports one model, while the rest of the network supports the competing one (see Figure 2(b) for an example). This indicates that confirmation bias has an important role in our model, facilitating polarization of opinions, since agents are influenced little by peers with conflicting opinions. The effect is larger as lambda increases, i.e. as under-reaction is growing, confirmation bias becomes more effective in generating polarization. This seems to indicate that a society that changes opinion slowly is more susceptible to polarisation due to confirmation bias.
@@
Another basic property of the model is that, while the model reaches consensus generally very fast, the time required to reach the stable state is increasing in $\lambda$ (as shown in Figure 3). In particular, as $\lambda$ approaches 1, nodes take a larger number of iterations to reach a stable configuration, because each of them is less sensitive to new information and under-reacts. The relation between the convergence time and $\phi$ is, however, not so linear. For low $\lambda$ values, large $\phi$ seems to slow down convergence, i.e. if the individuals do not under-react, then confirmation bias slows down convergence as society polarises. However, when $\lambda$ is large, the opposite effect is visible: since individuals under-react, polarisation is faster as confirmation bias increases: they are not attracted by peers from the opposite opinion groups and so the simulations ends faster.
@@
In Figure 4 we show the number of clusters and average subjective probabilities at the steady state. While in the baseline simulation outcomes were balanced between the two competing models (thus resulting in average subjective probabilities around 0.5 across independent runs, see Figure 1), in the one-lobbyist scenario average probabilities are significantly higher over the whole parameter space, indicating that generally the lobbyist is able to steer part of the network towards supporting its own preferred model. Furthermore, the effectiveness of the lobbyist is not globally homogeneous over the parameter space: in fact, the average probabilities are around $0.99$ in the bottom-left corner region (indicating that, with these configurations, the lobbyist is successful in influencing the network almost always), but tend to be lower as we go farther off said region. Moreover, we see that the lobbyist, while managing to successfully increase the average subjective probability distributions, is unable to eliminate or reduce clustering in the bottom-right corner of the map (Figure 4). If anything, in this region of the map an increase in polarization emerges, as the effective number of clusters increases, i.e. the size of the second cluster increases.
@@
These results indicate that a society with medium-high levels of under-reaction and low confirmation bias (dark blue area in Figure 4(b)) are easily influenced by a lobbyist, while when confirmation bias is strong or the population is more dynamic (low under-reaction), the peer-effect can contrast the lobbyist. This can generate consensus on the opposite opinion in a minority of simulations, when confirmation bias is low, or increase polarisation when confirmation bias is high. There appears to be a type of phase transition at the boundary, where the lobbyist does not have full power any more.
@@
Figure 4: One-lobbyist scenario. The effective number of clusters and average subjective probability of the final opinion distribution is represented as a function of the strength of directional motivated reason $\phi$ and the degree of under-reaction $\lambda$ in presence of a lobbyist. The lobbyists supports the pessimistic model, has a budget B=10, [^0] B=10,\!000 to send its signals and can be active for a time horizon $T=100$.
To investigate further this phase transition, we performed budget sensitivity analysis, i.e. changed the budget of the lobbyist from 10,000 to 5,000, 20,000 and 40,000, allowing them to reach 10%, 40% and 80% of the population, respectively. In Figure 5, we show the resulting average opinions. It is very clear that the region in which the lobbyist dominates the network (where it can impose its model almost certainly) expands linearly with the amount of resources available. With B>40, [^0] B>40,\!000, the lobbyist influence is complete and global over the parameter space. Even outside said region, however, average subjective probabilities are higher on average the higher the budget allocated to the lobbyist, indicating that the lobbyist has still a non-negligible impact on the network behaviour regardless of the specific parameter configuration.
(a) B=5, [^0] B=5,\!000
(b) B=10, [^0] B=10,\!000
(c) B=20, [^0] B=20,\!000
(d) B=40, [^0] B=40,\!000
@@
In this configuration, average subjective probabilities across simulations seem to balance out, just as in the baseline scenario, as it is shown in Figure 6(b). In fact, as the two lobbyists are identical except for the specific model supported, their competing influence seems to cancel out, and the average subjective probabilities gravitates around $\bar{p}=0.5$, regardless of the parametrization.
@@
Figure 6: The two-lobbyists scenario. The average number of iterations and average subjective probabilities of the final opinion distribution are represented as a function of strength of directional motivated reason $\phi$ and the degree of under-reaction $\lambda$. The lobbyists supports opposing models (one the optimistic model and the other the pessimistic one). Both of them are endowed with a budget B=10, [^0] B=10,\!000 to send its signals to the agents, i.e. can reach out an average of $20\%$ of the agents in each time steps and can be active for a time horizon $T=100$.
@@
The main difference emerging in this case in comparison to the baseline scenario is that competition between lobbyists over influence on the network results in an increase in the number of iterations needed to reach the stable state, as shown in Figure 6(a).
@@
In other words, the joint activity of the two lobbyists prevents the network from reaching an equilibrium as long as the lobbyists send signals; this is particularly true in the region that, in the previous scenario, was associated with higher lobbyist efficacy, strengthening the impression that, for that subset of configurations, it is easier for the external agent to project its influence on the network. Figure 6(c) represents the evolution of a typical run of the model in one of those configurations. Differently from what happens in the baseline scenario, where the network converges to a stable state relatively quickly, the system undergoes persistent fluctuations for as long as the two lobbyists are active on the network. Shortly after they stop emitting signals (beyond 100 iterations), however, the network quickly converges to a consensus, in a similar dynamics to what we have seen in the baseline scenario. It is worth noticing that while oscillations with stubborn influences also appear in linear models, here they emerge from motivated, signal-dependent responsiveness combined with time-staged strategic sending, yielding the regime patterns documented in Figs. 6-7.
@@
In this last scenario, we try to go a step further and we perform a sensitivity analysis on the lobbyists’ strategic horizon $T$ with the aim of investigating how this parameter affects the model’s properties. Specifically, we run several different batches of simulations by increasing the parameter $T$ to 1,000 and 2,000. At the same time, we adjust the lobbyists’ budget $B$ accordingly, in order to maintain the same rate of signals per time-period of 20%, thus keeping the results comparable across specifications. As shown in Figure 7, with a longer time horizon generally average opinion $\bar{p}$ remains about 0.5 in most configurations, suggesting that a longer horizon by the lobbyists does not radically change the outcome of the simulations. The average number of iterations, instead, tends to increase linearly with $T$ in the bottom-left corner of the parameter space, the region where the lobbyists are most effective , while it does not seem to change much in the other regions, that converge very fast due to peer effects.
(a) Average subjective probabilities for T=1, [^0] T=1,\!000
(b) Average subjective probabilities for T=2, [^0] T=2,\!000
(c) Average number of iterations for T=1, [^0] T=1,\!000
(d) Average number of iterations for T=2, [^0] T=2,\!000
@@
Comparing Lobbyists’ performance across different timing strategies
As a final exercise within the two-lobbyists scenario, we set up an experiment with the aim to investigate whether timing in the lobbyist strategies has a relevant impact on the outcome of the simulations. While insofar lobbyists have been assumed to spread out their signals uniformly over time, now we define two additional possible strategies, which we call "frontloading", and "backloading". In lobbying, frontloading means concentrating influence efforts at the earliest stages of the process. In the model, this is operationalized by the adoption, by the lobbyist agent, of a strategy matrix in which variable indicators for signals are concentrated in the first 20 (out of 100) iteration periods, in which the lobbyist exhausts its budget. Conversely, backloading implies holding back resources until later stages, in order to concentrate the effort closer to the steady state. This means, in our model, that the lobbyist adopts a strategy matrix in which variable indicators for signals are concentrated in the last 20 iteration periods.
For our experiment, we set up three different scenarios: in Scenario 1, both the pessimist and the optimist lobbyists adopt a frontloading strategy; in Scenario 2, the pessimist lobbyist adopts a frontloading strategy, while the optimist sticks to a random uniform strategy; in Scenario 3, the pessimist lobbyist adopts a backloading strategy, while the optimist keeps using a random uniform strategy. We run 100 Monte Carlo simulations for each parameter configuration and for each Scenario to assess the performance of the lobbyists under these different conditions.
Results are presented in Figure 8 (see also Figures S3 and S4 in Supplementary material). Expectedly, in the completely symmetric Scenario 1 (Figure 8(a)), average opinion outcomes are balanced across the parameter space, with neither lobbyist prevailing in the steady state across Monte Carlo simulations. The picture changes starkly in Scenario 2 (Figure 8(b)). In this case, as earlier in the model, we see two distinct regions emerge in the parameter space. In the region characterized by lower under-reaction (or high confirmation bias), where beliefs by nodes are more flexible and the peer effect is stronger, the frontloading (pessimist) lobbyist prevails over his opponent in almost all of the simulations. Conversely, the frontloading strategy seems to be defeated by the random uniform optimist lobbyist when under-reaction is higher and motivated reasoning component is lower. In contrast, in Scenario 3 (Fig 8(c)), the situation is reversed: the pessimist lobbyist (now backloading) dominates in the high-$\lambda$, low-$\phi$ configurations, but is dominated by the optimist elsewhere.
This behaviour can be readily interpreted by considering the core mechanisms driving the evolution of the network over time in our model. Frontloading is particularly effective when agents are highly responsive to new information and early signals can trigger reinforcing dynamics, such as cascades or clustering effects. By influencing opinions early, lobbyists can increase the likelihood that the system evolves toward their preferred equilibrium and can potentially reduce the need for later interventions.
Backloading, by contrast, targets the final opinion distribution directly, which is often the variable of interest for the lobbyist. Backloading is advantageous when agents under-react to signals or when early fluctuations are likely to reverse over time. By concentrating resources late, lobbyists maximize the impact of each signal on the final state, minimizing wasted influence and counteracting opposing signals. Backloading can be particularly effective in competitive environments where multiple lobbyists attempt to steer opinions, as late interventions can override prior influence and shift the final outcome.
Interestingly, the regions of parameter space where time-dependent strategies outperform the uniform strategy roughly match the regions in the single-lobbyist scenario where lobbyist influence is stronger (backloading prevails) and where peer effects are stronger (frontloading prevails).
(a) Scenario 1
(b) Scenario 2
(c) Scenario 3
Figure 8: Average steady-state subjective probabilities under different lobbying timing strategies. Panel (a) shows outcomes when both lobbyists adopt a frontloading strategy. Panel (b) depicts the case where the pessimist lobbyist frontloads while the optimist follows a uniform strategy. Panel (c) illustrates the reverse scenario with the pessimist backloading. Values are averaged over 100 independent runs for each setting. Simulations are performed on a fully connected network of $N=500$ agents and endowing each lobbyist with a budget $B=10,000$.
@@ Discussion
In this work, we introduce a model of opinion formation in which agents update beliefs via a Bayesian-like rule in a social-learning setting, incorporating cognitive biases such as under-reaction and directionally motivated reasoning (confirmation bias). The model includes lobbyists that are characterised by strategies (a set of agents to target at each iteration) and are budget constrained. In line with our scope, we study lobbying as an opinion-influence mechanism; the policy stage is exogenous and can be connected via a simple mapping without altering the core dynamics. After having provided some game theoretic examples, we studied the model numerically and observed the emergence of two polarised clusters when confirmation bias is large, even in the absence of lobbying. One lobbyist with an uniform strategy can attract the entire population to the opinion it promotes if the budget is large enough or if the individuals have low confirmation bias and medium under-reaction levels. Otherwise, a polarised minority continues to coexist with the cluster that agrees with the lobbyist. In the presence of two opposing lobbyists, the final steady states obtained are similar to those without lobbying, just that convergence happens only after the lobbyists stop influencing the network. Our oscillations are thus mechanistically distinct from those in linear stubborn-agent frameworks [1]: they require (i) nonlinearity in responsiveness (equation (3)) and (ii) endogenous, budget-constrained sending that injects temporally structured shocks (equations (4)-(5)). As long as they are active, the population continues to oscillate between different opinions without converging. This suggests that continuous lobbying can preclude consensus or convergence.
@@
Lobbyists are an important element of novelty in our model, since they are represented as more than an external field [12, 28, 42, 41, 36] or zealot/stubborn agent[32, 1]. Specifically, their effect on other agents can vary in time (at every iteration a subset of agents is reached) and is limited by a budget. This allows to model complex strategies and scenarios without changing the baseline model definition. We have shown that with this definition, one lobbyist attracts the complete population only in a certain area of the parameter space, a lobbyist influence area, where the population has low confirmation bias and medium levels of under-reaction. Polarisation appears as the population is more flexible (low under-reaction) or has more confirmation bias, where the peer effect can counteract the lobbyist influence. This polarising effect has also been seen in other models where extreme external information can cause minority extremist groups [42, 41]. A phase transition seems to appear at the boundary between these two regimes, i.e. lobbyist-influence area versus peer-effect area. When two symmetric opposing lobbyists are present, a rich behaviour is observed. During the active time horizon of the lobbyists, inside the lobbyist influence area, we observe fluctuations of opinions between the two extreme opinions, with convergence as soon as the lobbyists become silent. Finally, our analysis of alternative timing strategies shows that the distribution of signals is not neutral in general: frontloading can be highly effective when agents are responsive and cascades amplify early interventions, while backloading dominates when agents under-react and late interventions can directly shape the final steady state. This demonstrates that the temporal structure of lobbying is itself an additional strategic dimension of influence, and that its effectiveness depends to some degree on the cognitive bias of the population.
Future work will explore more complex lobbying strategies extracted from real data[19]. We will examine how these strategies perform across the two parameter regimes and run additional simulations on more realistic social networks (e.g., scale-free graphs) that better approximate real cases. We also plan to integrate the model with statistical model-checking environments, such as MultiVeStA[46, 40], to gain further insight into its properties and dynamics. This paper is a first step toward the broader goal of fully characterizing the complex strategic-interaction environment generated by rational lobbyists seeking to attract consensus on their positions.